a car is traveling on a road perpendicular to a railroad track. when the car is 30 meters from the crossing, new collision detector warns the driver that there is a train 50 meters from the car and heading toward the same crossing. How far is the train from the crossing?
Answer by richwmiller(17219)

a^2+b^2=c^2
a=30
c=50
30^2+b2=50^2
900+b^2=2500
b^2=1600
b=40 meters
this is a classic 3,4,5 right triangle